OpenStudy (anonymous):

Explain and help please! 29 + (8 + 17) Which of the following expressions is equivalent to the expression above? A. (29 + 8) + (29 + 17) B. (29 + 8) × (29 + 17) C. 29 + (17 + 8) D. 29 × 8 + 29 × 17

OpenStudy (anonymous):

See in addition (a+b)+c = a+(b+c)

OpenStudy (anonymous):

for example let a = 2, b = 3 and c = 4 then, (a+b)+c is (2+3)+4 = (5)+4 = 9 and similarly, a+(b+c) = 2+(3+4) = 2+7 = 9 So we can say that a+(b+c) = (a+b)+c

OpenStudy (anonymous):

Oh ok I understand now! Thanks I mean it!